# Newton’s method for nonlinear mappings into vector bundles

### v1

@@ 7.2. Variational problems on manifolds
which defines a covector back-transport $V_{x}^{*}\in\Gamma(\mathcal{L}(T^{*}\mathcal{X},T_{x}^{*}\mathcal{X}))$ on the cotangent bundle.
By using the natural inclusion $\varphi$ and differentiating ${\langle{V_{x}^{*}}\rangle}$, we obtain the dual connection map at $v^{*}\in T^{*}\mathcal{X}$ with $x=\pi^{*}(v^{*})\in\mathcal{X}$:
$Q_{v^{*}}^{*}:={\langle{V^{*}_{x}}\rangle}^{\prime}(v^{*}):T_{v^{*}}(T^{*}\mathcal{X})\rightarrow T_{x}^{*}\mathcal{X}.$
Combining (26) and (28) we obtain in trivializations
$Q^{*}_{v^{*}}\delta v^{*}\sim Q^{*}_{v^{*},\tau}(\delta x,\delta v^{*}_{\tau})=\delta v^{*}_{\tau}+v^{*}_{\tau}\circ R^{\prime\prime}_{x,\tau}(0_{x})(\delta x,\cdot).$ (30)
Embedded submanifolds.
Consider an embedded submanifold $\mathcal{X}$ of a Hilbert space $H$. We can use the embedding $E_{H}:T\mathcal{X}\to H$ with $E_{H}(x):T_{x}\mathcal{X}\to H$ and the orthogonal projection $P(\xi):H\to T_{\xi}\mathcal{X}$ on the tangent space to define a vector transport locally by
$V_{x}(\xi):=P(\xi)E_{H}(x)\in L(T_{x}\mathcal{X},T_{\xi}\mathcal{X}).$
A covector back-transport is then given by taking pointwise adjoints $V_{x}^{*}(\xi):=(V_{x}(\xi))^{*}$, i.e.
$V^{*}_{x}(\xi)\ell=\ell\circ V_{x}(\xi)=\ell\circ P(\xi)\circ E_{H}(x)\quad\forall\ell\in T_{\xi}^{*}\mathcal{X}.$
Consider a covector field $\ell_{H}:\mathcal{X}\to H^{*}$ and its restriction $\ell\in T^{*}\mathcal{X}$, given by $\ell(\xi):=\ell_{H}(\xi)\circ E_{H}(\xi)$. Differentiating ${\langle{V_{x}^{*}}\rangle}(\ell)(\xi)=\ell_{H}(\xi)\circ E_{H}(\xi)\circ V_{x}(\xi)$ with respect to $\xi$ at $\xi=x$ yields
${\langle{V_{x}^{*}}\rangle}^{\prime}(\ell(x))\ell^{\prime}(x)\delta x$ $=\ell_{H}^{\prime}(x)\delta x+\ell_{H}(x)(E_{H}\circ V_{x})^{\prime}(x)\delta x$
$=\ell_{H}^{\prime}(x)\delta x+\ell_{H}(x)(E_{H}\circ P)^{\prime}(x)E_{H}(x)\delta x,$
or, suppressing the embeddings notationally:
$Q^{*}_{\ell(x)}\ell^{\prime}(x)\delta x={\langle{V_{x}^{*}}\rangle}^{\prime}(\ell(x))\ell^{\prime}(x)\delta x=\ell_{H}^{\prime}(x)\delta x+\ell_{H}(x)P^{\prime}(x)\delta x\in T^{*}_{x}\mathcal{X}.$ (31)
Vice versa, if $\ell\in T^{*}\mathcal{X}$ is given, we may extend $\ell$ to $\ell_{H}:\mathcal{X}\to H^{*}$, for example by specifying $\ell_{H}(x)$ on $T_{x}\mathcal{X}^{\perp}$. Then by its intrinsic definition $Q^{*}_{\ell(x)}\ell^{\prime}(x)\delta x$ is independent of the extension, and thus also (31) is independent of the extension.
If $\mathcal{X}$ is given by a constraint of the form $\mathcal{X}=\{x\in H:c(x)=0\}$, where $c:H\to R$ is a differentiable mapping between Hilbert spaces, such that $c^{\prime}(x):H\to R$ is surjective, then we can find a Lagrangian multiplier $\lambda(x)\in R^{*}$, such that $(\ell_{H}(x)+\lambda(x)c^{\prime}(x))w=0$ for all $w\in T_{x}\mathcal{X}^{\perp}$. Using $\mathrm{ker}\,c^{\prime}(x)=T_{x}\mathcal{X}$ this relation can be written as
$0=(\ell_{H}(x)+\lambda(x)c^{\prime}(x))(Id_{H}-P(x))v=\ell_{H}(x)(Id_{H}-P(x))v+\lambda(x)c^{\prime}(x)v\quad\forall v\in H.$
$0=(\ell_{H}^{\prime}(x)\delta x(Id_{H}-P(x))-\ell_{H}(x)P^{\prime}(x)\delta x+\lambda^{\prime}(x)\delta x\,c^{\prime}(x)+\lambda(x)c^{\prime\prime}(x)\delta x)v\quad\forall v\in H$
$\lambda(x)c^{\prime\prime}(x)(\delta x,v)=\ell_{H}(x)P^{\prime}(x)\delta x\,v\quad\forall v\in T_{x}\mathcal{X}.$
Hence, (31) reads in this case:
$Q^{*}_{\ell(x)}\ell^{\prime}(x)\delta x=\ell_{H}^{\prime}(x)\delta x+\lambda(x)c^{\prime\prime}(x)\delta x\in T^{*}_{x}\mathcal{X}.$ (32)
@@
$Q_{f^{\prime}(x)}^{*}\circ f^{\prime\prime}(x)\delta x+f^{\prime}(x)=0_{x}^{*}.$ (33)
Since this is an equation in a dual space, $\delta x$ has to fulfill
@@
$*.\min_{\delta x\in T_{x}\mathcal{X}}\;(f\circ R_{x})^{\prime}(0_{x})\delta x+\frac{1}{2}(f\circ R_{x})^{\prime\prime}(0_{x})(\delta x,\delta x)\quad\Leftrightarrow\quad(f\circ R_{x})^{\prime\prime}(0_{x})\delta x+(f\circ R_{x})^{\prime}(0_{x})=0_{x}^{*}.$ (34)
@@
Thus, the Newton direction computed by (33) coincides with the Newton-SQP direction computed by (34) if $(f\circ R_{x})^{\prime\prime}(0_{x})$ is positive definite.
@@
Next, we have for $\delta x\in T_{x}\mathcal{X}$
$(f\circ R_{x})^{\prime\prime}(0_{x})(\delta x,v)=\frac{d}{d\xi}\left(f^{\prime}(R_{x}(\xi))R^{\prime}_{x}(\xi)\delta x\right)|_{\xi=0_{x}}v.$
Consider the trivializations $\tau_{T\mathcal{X}}:T\mathcal{X}\to\mathcal{X}\times\mathbb{X}$ and $\tau_{x}:\mathcal{X}\times T_{x}\mathcal{X}\to\mathcal{X}\times\mathbb{X}$. Using (1), we compute for $\xi\in T_{x}\mathcal{X}$:
$f^{\prime}_{\tau}(R_{x}(\xi))R_{x,\tau}^{\prime}(\xi)\delta x\circ\tau_{x}=(f^{\prime}(R_{x}(\xi))\circ\tau_{T\mathcal{X}}^{-1})\circ(\tau_{T\mathcal{X}}\circ R_{x}^{\prime}(\xi)\delta x\circ\tau_{x}^{-1})\circ\tau_{x}=f^{\prime}(R_{x}(\xi))R_{x}^{\prime}(\xi)\delta x.$
On the one hand, computing the derivative at $\xi=0_{x}$ yields, with $(R^{\prime}_{x,\tau}(0_{x})\circ\tau_{x})(v)=\tau_{x}(v)=v_{\tau}$ for $v\in T_{x}\mathcal{X}$:
$\frac{d}{d\xi}\left(f^{\prime}_{\tau}(R_{x}(\xi))R_{x,\tau}^{\prime}(\xi)\delta x\circ\tau_{x}\right)|_{\xi=0_{x}}v=f^{\prime\prime}_{\tau}(x)(\delta x,v_{\tau})+f^{\prime}_{\tau}(x)R^{\prime\prime}_{x,\tau}(0_{x})(\delta x,v_{\tau}).$
On the other hand, in trivializations we get with (30), using $v^{*}=f^{\prime}(x)$ and $\delta v^{*}\sim(\delta x,f_{\tau}^{\prime\prime}(x)\delta x)$:
$(Q_{f^{\prime}(x)}^{*}\circ f^{\prime\prime}(x)\delta x)v=f_{\tau}^{\prime\prime}(x)(\delta x,v_{\tau})+f_{\tau}^{\prime}(x)\circ R_{x,\tau}^{\prime\prime}(0_{x})(\delta x,v_{\tau}).$
This yields
$(f\circ R_{x})^{\prime\prime}(0_{x})\delta x=Q_{f^{\prime}(x)}^{*}\circ f^{\prime\prime}(x)\delta x\text{ in }T_{x}^{*}\mathcal{X}.$
Since $\delta x$ is an arbitrary tangent vector it holds that
Thus, the Newton equation (33) and the optimality constraint (34) coincide. ∎
@@
Our affine covariant damping strategy is well suited to find stationary points of $f$. If the computation of a local minimizer of $f$ is desired, different strategies from nonlinear optimization should be employed, such as linesearch or trust-region methods, that enforce descent in $f$ during the iteration.
If $\mathcal{X}$ is embedded into a Hilbert space $H$, $f_{H}:H\to\mathbb{R}$ and $f=f_{H}|_{\mathcal{X}}$, we may consider the constrained problem
@@
A stationary point is then given as $x\in\mathcal{X}$, such that $f^{\prime}(x)v=f_{H}^{\prime}(x)v=0$ for all $v\in T_{x}\mathcal{X}$. Then in view of (31) we obtain for the connection, induced by the orthogonal projection:
@@
In the case $\mathcal{X}=\{x\in H:c(x)=0\}$ this relation becomes
$Q^{*}_{f^{\prime}(x)}f^{\prime\prime}(x)\delta x=f_{H}^{\prime\prime}(x)\delta x+\lambda(x)c^{\prime\prime}(x)\delta x.$
where $(f^{\prime}_{H}(x)+\lambda(x)c^{\prime}(x))v=0$ for all $v\in T_{x}\mathcal{X}$ due to (32). Hence, in this case we retain the well known Lagrange-Newton step, used in constrained optimization.
@@ 8. Conclusion and future research
In our study of Newton’s method for a mapping $F:\mathcal{X}\to\mathcal{E}$ from a manifold into a vector bundle we have found a number of structural insights. The most basic distinction from the classical case is the need for a connection on $\mathcal{E}$ to render the Newton equation well defined. Together with a geometric version of Newton differentiability this already allows a local convergence theory of Newton’s method. A Banach type version of a Riemannian metric can be used as a flexible framework to formulate superlinear convergence.
For the development of further algorithmic strategies, like monitoring local convergence or globalization, vector back-transports are required. They make it possible to compare residuals from different fibres and to compute simplified Newton steps. We propose an affine covariant globalization scheme, which works purely with quantities that can be computed in terms of the domain $\mathcal{X}$. In the global regime, where residuals are not small, it is necessary that the employed connection is consistent with the vector back-transport in order to guarantee that the Newton direction is tangential to the algebraic Newton path. If $\mathcal{E}$ is the tangent bundle $T\mathcal{X}$ or the cotangent bundle $T^{*}\mathcal{X}$, then vector back-transports and connections can be derived from retractions on $\mathcal{X}$ by differentiation. In this way an algorithmic quantity on $\mathcal{X}$, a retraction, induces some geometric structure on $\mathcal{X}$, namely a connection.

### v2

@@ 7.3. Critical points of functionals
which defines a covector back-transport $V_{x}^{*}\in\Gamma(\mathcal{L}(T^{*}\mathcal{X},T_{x}^{*}\mathcal{X}))$ on the cotangent bundle. By differentiating ${\langle{V_{x}^{*}}\rangle}$, we obtain the dual connection map at $\ell\in T_{x}^{*}\mathcal{X}$:
$Q_{\ell}^{*}:={\langle{V^{*}_{x}}\rangle}^{\prime}(\ell):T_{\ell}(T^{*}\mathcal{X})\rightarrow T_{x}^{*}\mathcal{X}.$
Combining (36) and (40) we obtain in trivializations:
$Q^{*}_{\ell}\delta\ell\sim Q^{*}_{\ell,\tau}(\delta x,\delta\ell_{\tau})=\delta\ell_{\tau}+\ell_{\tau}\circ R^{\prime\prime}_{x,\tau}(0_{x})(\delta x,\cdot).$ (45)
@@
$\min_{\delta x\in T_{x}\mathcal{X}}\;\frac{1}{2}(f\circ R_{x})^{\prime\prime}(0_{x})(\delta x,\delta x)+(f\circ R_{x})^{\prime}(0_{x})\delta x\quad\Leftrightarrow\quad(f\circ R_{x})^{\prime\prime}(0_{x})\delta x+(f\circ R_{x})^{\prime}(0_{x})=0_{x}^{*}.$ (47)
@@
Thus, if $(f\circ R_{x})^{\prime\prime}(0_{x})$ is positive definite, the Newton direction computed by (46) coincides with the Newton-SQP direction computed by (47), if any of them exists.
@@
To obtain the second equality we choose $\delta x\in T_{x}\mathcal{X}$ and apply Lemma 7.4 to $F(x)=f^{\prime}(x)$, with $y(x)=x$, $V_{x}(\xi)=R^{\prime}_{x}(R_{x}^{-1}(\xi))$. Setting $\xi=R_{x}(w)$ for $w\in T_{x}\mathcal{X}$ and using $(R_{x}^{-1})^{\prime}(x)=Id_{T_{x}\mathcal{X}}$ we can compute:
$(f\circ R_{x})^{\prime\prime}(0_{x})(v,\delta x)$ $=\frac{d}{dw}\Big{(}f^{\prime}(R_{x}(w))R^{\prime}_{x}(w)v\Big{)}\Big{|}_{w=0_{x}}\delta x$
[…]
Thus, the Newton equation (46) and the optimality condition (47) coincide. ∎
@@
Our affine covariant damping strategy is well suited to finding stationary points of $f$. If the computation of a local minimizer of $f$ is desired, different strategies from nonlinear optimization should be employed, such as line-search or trust-region methods that enforce descent in $f$ during the iteration.
7.4. Embedded vector bundles
Consider a vector bundle $\mathcal{E}$ that is embedded into a Hilbert space $H$. Then we can use the embedding $\iota_{H}:\mathcal{E}\to H$ with $\iota_{H}(y):E_{y}\to H$ and the orthogonal projection $P(y):H\to E_{y}$ onto the fibre to define a vector back-transport locally by
$V_{y}^{-1}(\eta):=P(y)\iota_{H}(\eta)\in L(E_{\eta},E_{y}).$
Analogously, we can define a vector transport locally by
$V_{y}(\eta):=P(\eta)\iota_{H}(y)\in L(E_{y},E_{\eta}).$
As seen before, we may obtain a consistent connection map by differentiation. Let us now consider a vector bundle $p:\mathcal{E}\to\mathcal{Y}$ and its dual bundle $p^{*}:\mathcal{E}^{*}\to\mathcal{Y}$. By our natural inclusion $\varphi$, we obtain the $C^{1}$-mapping ${\langle{V_{y}^{-1}}\rangle}:\mathcal{E}\to E_{y}$ with ${\langle{V_{y}^{-1}}\rangle}(e)({\eta})=P(y)(\iota_{H}e)({\eta})$ for $e\in\mathcal{E}$ with $p(e)=y$. Differentiating this with respect to ${\eta}$ at ${\eta}=y$, we obtain
$Q_{e(y)}e^{\prime}(y)\delta y={\langle{V_{y}^{-1}}\rangle}^{\prime}(e(y))e^{\prime}(y)\delta y=P(y)(\iota_{H}e)^{\prime}(y)\delta y\in E_{y}.$ (48)
For a vector field $\nu:\mathcal{X}\to T\mathcal{X}$, this projection based vector back-transport can be used as an alternative to the retraction based transport (35). Using (48) and $\pi(\nu(x))=x$ for $x\in\mathcal{X}$, we obtain
$Q_{\nu(x)}\nu^{\prime}(x)\delta x={\langle{V_{x}^{-1}}\rangle}^{\prime}(\nu(x))\nu^{\prime}(x)\delta x=P(x)(\iota_{H}\nu)^{\prime}(x)\delta x\in T_{x}\mathcal{X},$
which is consistent with the orthogonal projection. The Newton equation reads in this case
$P(x)(\iota_{H}\nu)^{\prime}(x)\delta x+\nu(x)=0_{x},\quad\delta x\in T_{x}\mathcal{X}.$
Remark 7.7.
The covariant derivative $Q\circ\nu^{\prime}$ of a vector field $\nu$ w.r.t. the connection map $Q$ can therefore be determined by orthogonally projecting the derivatives of the embedded vector field onto $T\mathcal{X}$. The connection which belongs to this covariant derivative is sometimes called tangential connection and defines the Levi-Civita connection on $T\mathcal{X}$, see e.g. [15, Chap. 5].
Dually, consider $\ell_{H}:\mathcal{Y}\to H^{*}$ and its restriction $\ell\in\mathcal{E}^{*}$, given by $\ell({\eta}):=\ell_{H}({\eta})\circ\iota_{H}({\eta})$. A covector back-transport is given by taking pointwise adjoints $V_{y}^{*}({\eta}):=(V_{y}({\eta}))^{*}$, i.e.
$V^{*}_{y}({\eta})\ell=\ell\circ V_{y}({\eta})=\ell\circ P({\eta})\circ\iota_{H}(y)\quad\forall\ell\in E_{\eta}^{*}.$
Differentiating ${\langle{V_{y}^{*}}\rangle}(\ell)({\eta})=\ell_{H}({\eta})\circ\iota_{H}({\eta})\circ V_{y}({\eta})$ with respect to ${\eta}$ at ${\eta}=y=p^{*}(\ell)$ yields
${\langle{V_{y}^{*}}\rangle}^{\prime}(\ell(y))\ell^{\prime}(y)\delta y$ $=\ell_{H}^{\prime}(y)\delta y+\ell_{H}(y)(\iota_{H}\circ V_{y})^{\prime}(y)\delta y$
$=\ell_{H}^{\prime}(y)\delta y+\ell_{H}(y)(\iota_{H}\circ P)^{\prime}(y)\iota_{H}(y)\delta y,$
or, suppressing the embeddings notationally:
$Q^{*}_{\ell(y)}\ell^{\prime}(y)\delta y={\langle{V_{y}^{*}}\rangle}^{\prime}(\ell(y))\ell^{\prime}(y)\delta y=\ell_{H}^{\prime}(y)\delta y+\ell_{H}(y)P^{\prime}(y)\delta y\in E_{y}^{*}.$ (49)
Vice versa, if $\ell\in\mathcal{E}^{*}$ is given, we may extend $\ell$ to $\ell_{H}:\mathcal{Y}\to H^{*}$, for example by specifying $\ell_{H}(y)$ on $\iota_{H}(E_{y})^{\perp}$. Then by its intrinsic definition $Q^{*}_{\ell(y)}\ell^{\prime}(y)\delta y$ is independent of the extension, and thus also (49) is independent of the extension.
Constrained optimization.
For $f_{H}:H\to\mathbb{R}$ and $f=f_{H}|_{\mathcal{X}}$, we may apply our results to the constrained optimization problem
@@ 7.4. Embedded vector bundles
A stationary point is then given as $x\in\mathcal{X}$, such that $f^{\prime}(x)v=f_{H}^{\prime}(x)v=0$ for all $v\in T_{x}\mathcal{X}$. In view of (49) setting $\ell_{H}=f_{H}^{\prime}$ we obtain for the connection, induced by the orthogonal projection:
@@
In optimization $\mathcal{X}$ is often given by an equality constraint of the form $\mathcal{X}=\{x\in H:c(x)=0\}$, where $c:H\to R$ is a differentiable mapping between Hilbert spaces, such that $c^{\prime}(x):H\to R$ is surjective, with $\mathrm{ker}\,c^{\prime}(x)=T_{x}\mathcal{X}$. Then the restriction $c^{\prime}(x):T_{x}\mathcal{X}^{\perp}\to R$ is an isomorphism, and so is its adjoint. This means that there is a unique Lagrangian multiplier $\lambda(x)\in R^{*}$, satisfying
$(f^{\prime}_{H}(x)+\lambda(x)c^{\prime}(x))w=0\quad\forall w\in T_{x}\mathcal{X}^{\perp}.$
Using $\mathrm{ker}\,c^{\prime}(x)=T_{x}\mathcal{X}=\mathrm{ran}\,P(x)$ this relation can be written as
$0=(f^{\prime}_{H}(x)+\lambda(x)c^{\prime}(x))(Id_{H}-P(x))v=f^{\prime}_{H}(x)(Id_{H}-P(x))v+\lambda(x)c^{\prime}(x)v\quad\forall v\in H.$
$0=\Big{(}(f^{\prime\prime}_{H}(x)\delta x)(Id_{H}-P(x))-f^{\prime}_{H}(x)P^{\prime}(x)\delta x+(\lambda^{\prime}(x)\delta x)\,c^{\prime}(x)+\lambda(x)c^{\prime\prime}(x)\delta x\Big{)}v\quad\forall v\in H$
$\lambda(x)c^{\prime\prime}(x)(\delta x,v)=f^{\prime}_{H}(x)P^{\prime}(x)\delta x\,v\quad\forall v\in T_{x}\mathcal{X}.$
Hence, (49) reads in this case
$Q^{*}_{f^{\prime}(x)}f^{\prime\prime}(x)\delta x=f^{\prime\prime}_{H}(x)\delta x+\lambda(x)c^{\prime\prime}(x)\delta x\in T^{*}_{x}\mathcal{X}.$ (50)
which yields the well known Hessian of the Lagrangian. An SQP-step at $x\in\mathcal{X}$ can then be computed by the following well known linearly constrained problem:
$\min_{\delta x\in H}f_{H}^{\prime}(x)\delta x+\frac{1}{2}(f_{H}^{\prime\prime}(x)+\lambda(x)c^{\prime\prime}(x))(\delta x,\delta x)\quad\text{s.t.}\quad c^{\prime}(x)\delta x=0.$
@@ 8. Conclusion and future research
In our study of Newton’s method for a mapping $F:\mathcal{X}\to\mathcal{E}$ from a manifold into a vector bundle we have found a number of structural insights. The most basic distinction from the classical case is the need for retractions $R_{x}$ on $\mathcal{X}$ and connections $Q_{e}$ on $\mathcal{E}$ to render the Newton steps well defined. Together with a geometric version of Newton-differentiability this already allows a local convergence theory of Newton’s method. A Banach type version of a Riemannian metric can be used as a flexible framework to formulate superlinear convergence.
For the development of further algorithmic strategies, like monitoring local convergence or globalization, vector back-transports $V_{y}^{-1}$ on $\mathcal{E}$ are required. They make it possible to compare residuals from different fibres and to compute simplified Newton steps. We propose an affine covariant globalization scheme, which works purely with quantities that can be computed in terms of the domain $\mathcal{X}$. In the global regime, where residuals are not small, it is necessary that the employed connection is consistent with the vector back-transport in order to guarantee that the Newton direction is tangential to the algebraic Newton path. If $\mathcal{E}$ is the tangent bundle $T\mathcal{X}$ or the cotangent bundle $T^{*}\mathcal{X}$, then vector back-transports and connections can be derived from retractions on $\mathcal{X}$ by differentiation. In this way an algorithmic quantity on $\mathcal{X}$, a retraction, induces some geometric structure on $\mathcal{X}$, namely a connection.